The distance between 100 and me is 28. What numbers can i be
Answer by stanbon(75887)

The distance between 100 and me is 28. What numbers can i be
----
Solve |x-100| = 28
---
x-100 = 28 or x-100 = -28
x = 128 or x = 72
==========================
cheers,
Stan H.
========